OpenStudy (anonymous):

Help to check my work: In a survey of 500 people, 200 indicated that they could be buying a major appliance within the next month, 150 indicated that they would buy a car, and 25 said that they would purchase both a major appliance and a car. How many will purchase neither. How many will purchase only a car?

OpenStudy (anonymous):

\[n(\Omega)\]: Total number of people.

OpenStudy (anonymous):

\[n(A)\]: number of people that will buy a major appliance.

OpenStudy (anonymous):

\[n(B)\]: number of people that will buy a car.

OpenStudy (anonymous):

\[n(C)\]: number of people that will buy both.

OpenStudy (anonymous):

\[C = A intersection B\]

OpenStudy (anonymous):

So, to get the howmany will purchase a car:

OpenStudy (anonymous):

\[n(A\\C)=n(A)-n(A\intersection B)\]

OpenStudy (anonymous):

So, \[n(A\C)=200 - 25 = 175\]

OpenStudy (turingtest):

I'm not sure about the formulas for this kind of thing but I get the same answer.

OpenStudy (mertsj):

150 will buy a car.

OpenStudy (turingtest):

oh no I don't, I get 125 for the 'neither' part

OpenStudy (mertsj):

The problem tells you that.

OpenStudy (anonymous):

And to get How many will purchase either: \[n((A \cup B)^C)=n(\Omega)-n(A\cup B)=n(\Omega) - n(A)-n(B)+n(A intersection B)\]

OpenStudy (mertsj):

Draw a Venn Diagram

OpenStudy (anonymous):

I haven't calculated yet, but I will use that formula \TuringTest

OpenStudy (anonymous):

I did it to get my formulas @Mertsj

OpenStudy (turingtest):

why are you adding the intersection in that formula? is that right above?

OpenStudy (anonymous):

Because : \[n(A\cup B) = n(A) + n(B) - n(A intersection B)\]

OpenStudy (anonymous):

I got 175 for the neither part =(

OpenStudy (turingtest):

oh, is it because we counted the people who will buy one or the other in with those who will buy both? then it makes sense

OpenStudy (turingtest):

ok then the number of people who will buy JUST a car is not 150

OpenStudy (anonymous):

yes, is say, How many will buy only

OpenStudy (anonymous):

a car

OpenStudy (anonymous):

So, do you think I am right?

OpenStudy (mertsj):

125 will buy only a car

OpenStudy (anonymous):

Why Mertsj?

OpenStudy (anonymous):

What is wrong with my formula?

OpenStudy (mertsj):

Because it says that 150 will buy a car and 25 will buy both a car and an appliance. 150-25 = 125

OpenStudy (anonymous):

Ahh yes! I made a mistake with the numbers. sorry

OpenStudy (anonymous):

but what do you think of the neither part?

OpenStudy (turingtest):

It also makes the neither part 150, yes?

OpenStudy (anonymous):

I guess I'm over complicating the things

OpenStudy (mertsj):

500 -[200+150-25]=500-325=175

OpenStudy (anonymous):

No I still get 175

OpenStudy (turingtest):

because 200 buy a car and 150 buy an appliance (forget the "both" it doesn't matter) 350 bought something so 500-350=150 did not

OpenStudy (turingtest):

I git the 200 and the 150 backwards, but it is the same idea

OpenStudy (mertsj):

Yes you are over complicating things and going around Robin Hood's barn to boot.

OpenStudy (anonymous):

But if you forget about the "both" part you're counting twice?

OpenStudy (mertsj):

No Turing. You are counting the 25 who bought both twice.

OpenStudy (turingtest):

oh, am I?

OpenStudy (anonymous):

What is the meaning of Robin Hood's barn to boot?

OpenStudy (mertsj):

175 bought only a MA. 125 bought only a car. 25 bought both. So 325 bought something.

OpenStudy (anonymous):

Yes, I think there is no reason to forget the both part.

OpenStudy (turingtest):

oh ok... my bad :(

OpenStudy (anonymous):

Anyway, thank you so mucho for the help guys =).

OpenStudy (mertsj):

yw